OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and has a slope of 2 in point-slope form

OpenStudy (sohailiftikhar):

as point slope form:_ Y=mx+c or y-y1=m(x-x1) you are given slope m=2 , x1=1 and y1=3 then y-3=2(x-1) 2x-2=y-3 2x-y+3-2=0 2x-y+1=0 that is the required equetion ...